# AlphaNet: Scaling Up Local Frame-based Neural Network Interatomic Potentials

### v3

@@ Abstract
Molecular dynamics simulations demand unprecedented accuracy and scalability to tackle grand challenges in energy materials, catalytic processes, and biomolecular design. To bridge this gap, we present AlphaNet, a local frame-based equivariant model that simultaneously advances computational efficiency and predictive precision for atomistic systems. By constructing equivariant local frames with learnable geometric transitions, AlphaNet encodes atomic environments with enhanced representational capacity, achieving state-of-the-art accuracy in energy and force predictions. Extensive benchmarks—spanning defected graphene, formate decomposition, inorganic bulks, and large-scale datasets (OC2M and Matbench Discovery) demonstrate its superior performance over existing neural network interatomic potentials while ensuring scalability across diverse system sizes. The synergy of accuracy, efficiency, and transferability positions AlphaNet as a transformative tool for simulating multiscale phenomena, from catalyst dynamics to energy storage interfaces, with direct implications for accelerating the discovery of functional materials and complex molecular systems. Our code and data is available at https://github.com/zmyybc/AlphaNet.
00footnotetext: These authors contributed equally to this work.00footnotetext: Corresponding authors. Emails: wangjiaao0720@utexas.edu, yuanqidu@cs.cornell.edu, duanchenru@gmail.com and haixiao@tsinghua.edu.cn.
@@ 1 Introduction
Molecular dynamics (MD) simulations have become essential for exploring and understanding complex phenomena in areas such as energy storage, catalysis, and biological systems (Richards et al., 2016; Boero et al., 1998; Lindorff-Larsen et al., 2011). While atomic forces necessary for these simulations can be derived from quantum mechanical approaches such as density functional theory (DFT), the computational cost associated with such first-principles methods severely limits their application to small-sized systems ($\sim{}10^{3}$ atoms) and short timescales ($\sim{}10^{1}$ ps). As a result, a wide spectrum of phenomena occurring over much larger scales remain inaccessible, even with the most powerful supercomputers.
@@
Classical force fields, which rely on predefined mathematical forms, offer a computationally efficient alternative, allowing simulations of larger systems over extended trajectories (Jones and Chapman, 1924; Cornell et al., 1995; MacKerell et al., 1998). However, the simplicity of these models often compromises their accuracy, leading to a trade-off between computational speed and the fidelity of the simulated dynamics (Iftimie et al., 2005).
@@
The advent of machine learning techniques has introduced a promising solution to this dilemma. By training models on data derived from first-principles calculations, neural network interatomic potentials (NNIPs) can potentially achieve the accuracy of first-principles methods while maintaining the computational efficiency of classical force fields. Unlike traditional models that rely on explicit functional forms to describe the bonded and non-bonded interactions, NNIPs are more flexible, learning to predict interactions based on the positions of atoms and their chemical identities.
@@
As interest in applying NNIPs to large-scale atomistic systems continues to grow, the field has seen a surge in the development of various innovative models. One example among the others, DPA-1 (Zhang et al., 2024a), is a large-scale pre-trained model with improved attention architecture over DeepPot (Wang et al., 2018). Similarly, JMP (Shoghi et al., 2024) leverages diverse molecular systems of different types as a joint pre-training model. MACE (Batatia et al., 2022), on the other hand, introduces higher-order message passing as a complete basis of many-body atomic interactions. Collectively, these models represent significant strides in the evolution of NNIP-driven atomistic simulations (Qu and Krishnapriyan, 2024; Liao et al., 2024; Gasteiger et al., 2021; Batzner et al., 2022; Schütt et al., 2018; Smith et al., 2017; Deng et al., 2023).
@@
Nevertheless, the balance between computational efficiency and accuracy is key to the practical application of these models. On one side, efficient yet less expressive models are more suitable for tasks that weigh computational cost more than accuracy. On the other side, less efficient yet expressive models are more suitable for tasks where accuracy is more important. Atomistic systems in 3D Euclidean space are invariant to Euclidean symmetries including rotation, translation and reflection. Thus, equivariant models are developed to respect these symmetries  (Satorras et al., 2021; Thomas et al., 2018), and most of these expressive models by date (Batzner et al., 2022; Batatia et al., 2022; Liao et al., 2024) are based on the spherical harmonics (Geiger and Smidt, 2022), but calculating the tensor product of irreducible representations imposes expensive computational overheads. While another branch of work achieves rotation equivariance through building equivariant frames that can be either local or global (Jumper et al., 2021; Du et al., 2022, 2023; Duval et al., 2023), and the main benefit of frame-based approaches is eliminating the necessity to involve the tensor product, thus greatly improving the computational efficiency.
@@
In this paper, we propose a local-frame-based equivariant atomistic model, named AlphaNet, for accurate yet highly efficient atomistic simulations. Building on the success of frame-based NNIPs in small atomistic systems, we introduce an additional rotary positional embedding to enhance the frame transition and temporal connection for multi-scale modeling. Extensive quantitative experiments demonstrate that AlphaNet excels at accuracy, efficiency and scalability compared with the state-of-the-art (SOTA) NNIP models on a variety of systems, ranging from defected graphenes, formate decomposition, zeolites, surface reactions and inorganic bulks.
@@ 2 Results
We systematically assess the capabilities of AlphaNet by examining its prediction accuracy, scalability with respect to model and dataset size, and computational efficiency in terms of inference speed and memory usage. Initially, we validate atomic-level prediction accuracy across five diverse chemical systems, simultaneously demonstrating AlphaNet’s scalability concerning dataset and system sizes. Subsequently, we quantify computational efficiency by benchmarking both the memory footprint and inference costs of the model. Comprehensive details regarding datasets, comparative methodologies, and hyperparameter configurations are provided in Appendix B.
2.1 AlphaNet achieving high accuracy across all datasets.
We first consider the formate decomposition dataset, which serves as a representative example of catalytic surface reactions, specifically focuses on the dehydrogenation of formate (HCOO* → H* + CO2) on a Cu $\langle 110\rangle$ surface. For this dataset, AlphaNet achieves a mean absolute error (MAE) of 45.5 meV/Åfor force and 0.23 meV/atom for energy, compared to NequIP (Batzner et al., 2022) 47.3 meV/Åand 0.50 meV/atom, respectively, as shown in Table 1. These results underscore AlphaNet’s ability to effectively capture the intricate nature of heterogeneous systems, which often involve both metallic and covalent bonding, along with complex charge transfer processes between the metal and adsorbed molecules. This high degree of accuracy highlights the model’s suitability for simulating catalytic reactions with multiple interaction types.
Another challenge for NNIP models is accurately modeling the sliding layer effect in materials like defected graphene, where inter-layer interactions are critical. On the defected graphene dataset, AlphaNet achieves a force MAE of 32.0 meV/Åand an energy MAE of 1.7 meV/atom, substantially outperforming NequIP, which records 60.2 meV/Åfor force MAE and 1.9 meV/atom for energy MAE. Furthermore, AlphaNet successfully reproduces the binding energy curve for AB-stacked bilayer graphene as calculated by PBE+MBD, without requiring explicit long-range dispersion interaction corrections. The deviation in the shallow sliding potential energy landscape is less than 0.4 meV/atom, illustrating AlphaNet’s capacity to handle systems with subtle interlayer forces and complex structural dynamics (as shown in Figure 1).
@@
Defected Graphene Force MAE (meV/Å) $\downarrow$ 60.2 32.0
Energy MAE (meV/atom) $\downarrow$ 1.9 1.7
@@
The third dataset we consider is the zeolite dataset, comprising 16 types of zeolites and a total of 800,000 configurations. Table 2 shows the results of Deep Pot (Wang et al., 2018) and AlphaNet across different zeolite configurations. Across nearly all tested configurations, AlphaNet significantly outperforms Deep Pot in both energy and force predictions. For instance, in the ABWopt configuration, AlphaNet achieves an Energy MAE of 41 meV, substantially lower than Deep Pot’s 90 meV. Similarly, AlphaNet records a Force MAE of 54 meV/Å, outperforming Deep Pot’s 90 meV/Å.
@@
Energy MAE $\downarrow$ 90 71 110 73 210 41 150 62
Force MAE $\downarrow$ 90 54 50 55 60 90 90 10
@@
Energy MAE $\downarrow$ 40 5.4 290 100 60 56 150 18
Force MAE $\downarrow$ 50 59 130 18 50 3.1 70 50
@@
Energy MAE $\downarrow$ 110 130 150 90 70 1.2 210 47
Force MAE $\downarrow$ 70 46 64 60 110 60 110 9.9
@@
Energy MAE $\downarrow$ 90 6.5 60 16 200 9.5 160 70
Force MAE $\downarrow$ 70 3.5 120 50 110 21 60 47
@@
The next challenge, we consider the Open Catalyst Project OC20 dataset (Chanussot et al., 2021), a comprehensive collection of crystal structures for training NNIP models in catalysis. We focus on the subset OC2M and the Structure-to-energy-and-force (S2EF) task. We compare the performance of our model against established benchmarks, including EquiformerV2 (Liao et al., 2024), EScAIP (Qu and Krishnapriyan, 2024), eSCN (Passaro and Zitnick, 2023), GemNet-OC (Gasteiger et al., 2022), SchNet (Schütt et al., 2018), and DimeNet++ (Gasteiger et al., 2020).
AlphaNet, trained on the 2M subset for 2,000,000 steps, achieves exceptional accuracy in energy prediction, with a mean absolute error (MAE) of 0.24 eV. This performance is comparable to larger models like EquiformerV2 and EScAIP, while SchNet and DimeNet++, trained on the full dataset, exhibit worse MAE values (>0.35 eV). However, AlphaNet trained on the 2M subset does not outperform larger models like EquiformerV2 and EScAIP in the force prediction task. This may be attributed to the conservativeness of the models, which ensures forces are derived as $-\nabla_{x}E(x)$.
@@
EScAIP-Small 0.782 5.634 0.712 0.939 0.038 0.783 45M EFSD
AlphaNet 0.799 4.863 0.743 0.933 0.041 0.745 16.2M EFSG
For the last challenge, we validate the effectiveness of AlphaNet on the well-established Matbench discovery WBM test set (Riebesell et al., 2024), AlphaNet demonstrates a compelling balance between accuracy and physical consistency on the WBM benchmark 4. With a 0.799 F1 score and 4.863 DAF score, it achieves near state-of-the-art classification performance while employing an EFSG-type conservative force field that rigorously enforces physical constraints. Notably, its compact architecture (16.2M parameters) outperforms the 31.2M-parameter eqV2 S DeNS model in regression tasks ($R^{2}=0.742$) with lower prediction error ($\text{MAE}=0.039$). This combination of efficiency and physics-aware design positions AlphaNet as a robust NNIP model for scalable materials modeling.
2.2 AlphaNet is scalable—effective scaling with both model, data and system sizes
In addition to maintaining high performance with large datasets and large systems, we validate the scaling of model size and dataset size on the zeolite dataset. We can observe the performance improves as the training data size increases from 50k to 800k samples, as shown in Figure 2. Specifically, the energy MAE decreases from 220 meV at 50k samples to 41 meV at 800k samples. Similarly, the force MAE decreases from 241 meV/Åat 5k samples to 54 meV/Åat 80k samples. Furthermore, deeper neural networks improve performance faster with the growth of the dataset size. This trend highlights the scalability of AlphaNet with increasing dataset size and model size.
@@ 2.2 AlphaNet is scalable—effective scaling with both model, data and system sizes
2.3 AlphaNet is efficient—achieves the best speed-accuracy trade-off
The primary advantage of NNIP models over density functional theory (DFT) calculated energy and force lies in their significantly faster inference speed. Although increasing the number of parameters in an NNIP model generally enhances its performance, this often comes at the cost of slower inference speed. To evaluate the speed of different NNIPs, we evaluate the average forward pass time across 2,000 zeolite structures with a batch size of 10. In addition, we test systems of varying sizes by expanding the unit cells. Although we aim to maintain a consistent number of parameters across models, slight variations exist due to different components in each model (detailed in Section B.4).
Figure 3: Comparisons of inference speed among various NNIP models. (a) Inference speed as a function of system size. The x-axis represents the number of atoms in the system, and the y-axis shows the average time required to predict energy and forces for systems in batches of 10, across a total of 200 batches on the Zeolite dataset. (b) Inference speed, F1 performance and model size of models accessible and compliant with the Matbench Discovery leaderboard. The speed is calculated as the number of energy evaluation per second.
As illustrated in Figure 3 (a), AlphaNet demonstrates a distinct speed advantage over other NNIP models, exhibiting a relatively slow increase in computational time as the size of the system increases. This indicates that AlphaNet’s speed scales efficiently with larger systems, making it a more suitable option for handling increasing system sizes.
Furthermore, we evaluated the inference speed of the pre-trained models on the Matbench Discovery leaderboard (Riebesell et al., 2024). To ensure fairness, we select models that are compliant, open-source, and accessible without additional requests. We evaluate the average speed of energy prediction and present in Figure 3 (b). The models included AlphaNet, ORB-v2 (Neumann et al., 2024), SevenNet (Park et al., 2024), GRACE (Bochkarev et al., 2024), MACE (Batatia et al., 2022), and CHGNet (Deng et al., 2023). Among these models, GRACE, developed based on TensorFlow, exhibits remarkable speed compared to the others, which are PyTorch-based. Notably, AlphaNet-MPtrj’s inference speed is comparable to models with 10 times fewer parameters (e.g., SevenNet-l3i5). This suggests that AlphaNet achieves competitive efficiency despite its larger parameter count, and it would be intriguing to explore its performance with fewer parameters in future work.

### v4

# AlphaNet: Scaling Up Local-frame-based Neural Network Interatomic Potentials

@@ Abstract
Molecular dynamics simulations demand an unprecedented combination of accuracy and scalability to tackle grand challenges in catalysis and materials design. To bridge this gap, we present AlphaNet, a local-frame-based equivariant model that simultaneously improves computational efficiency and predictive precision for interatomic interactions. By constructing equivariant local frames with learnable geometric transitions, AlphaNet encodes atomic environments with enhanced representational capacity, achieving state-of-the-art accuracy in energy and force predictions. Extensive benchmarks on large-scale datasets spanning molecular reactions, crystal stability, and surface catalysis (Matbench Discovery and OC2M) demonstrate its superior performance over existing neural network interatomic potentials while ensuring scalability across diverse system sizes with varying types of interatomic interactions. The synergy of accuracy, efficiency, and transferability positions AlphaNet as a transformative tool for modeling multiscale phenomena, decoding dynamics in catalysis and functional interfaces, with direct implications for accelerating the discovery of complex molecular systems and functional materials. Our code and data are available at https://github.com/zmyybc/AlphaNet.
@@ 1 Introduction
Molecular dynamics (MD) simulations have become essential for investigating and elucidating complex phenomena across diverse fields, including biology, catalysis, and energy engineering [21, 20, 1, 2, 27, 43, 9, 37]. While atomic forces necessary for MD simulations can be derived from quantum mechanical approaches such as density functional theory (DFT) [10], the computational cost associated with such first-principles methods severely limits their applications to small-sized systems ($\sim{}10^{3}$ atoms) and short timescales ($\sim{}10^{1}$ ps). Consequently, a wide spectrum of phenomena occurring over much larger scales remain inaccessible, even with the most powerful supercomputers.
@@
Classical force fields, which rely on predefined mathematical forms, offer a computationally efficient alternative, allowing simulations of larger systems over extended trajectories [29, 14, 38, 47, 55, 31]. However, the simplicity of these models often compromises their accuracy, necessitating a delicate trade-off between computational efficiency and the fidelity of the simulated dynamics, especially for reactive chemistry [55, 53, 28, 64].
@@
The advent of machine learning techniques has introduced a promising solution to this dilemma. By training models on data derived from first-principles calculations, neural network interatomic potentials (NNIPs) can potentially achieve the accuracy of first-principles methods while maintaining the computational efficiency of classical force fields [4, 5, 63, 6, 3, 51, 32]. Unlike traditional models that rely on explicit functional forms to describe the bonded and non-bonded interactions, NNIPs are more flexible, learning to predict interactions based on the positions of atoms and their chemical identities [7].
@@
As interest in applying NNIPs to large-scale atomistic systems continues to grow, the field has seen a surge in the development of various innovative models. One example among the others, DPA-1 [62], is a large-scale pre-trained model with improved attention architecture over DeepPot [56]. Similarly, JMP [49] leverages diverse molecular systems of different types as a joint pre-training model. MACE [4], on the other hand, introduces higher-order message passing as a complete basis of many-body atomic interactions. Collectively, these models represent significant strides in the evolution of NNIP-driven atomistic simulations [42, 36, 23, 5, 46, 50, 15, 13, 60].
@@
Nevertheless, the balance between computational efficiency and accuracy is key to practical applications of these models. On one side, efficient yet less expressive models are more suitable for tasks that weigh computational cost more than accuracy. On the other side, less efficient yet expressive models are more suitable for tasks where accuracy is more important. Atomistic systems in 3D Euclidean space are invariant to Euclidean symmetries including rotation, translation and reflection. Thus, equivariant models are developed to respect these symmetries  [45, 54], and most of these expressive models by date [5, 4, 36] are based on the spherical harmonics [25], but calculating the tensor product of irreducible representations imposes expensive computational overheads. While another branch of work achieves rotation equivariance through building equivariant frames that can be either local or global [30, 16, 17, 19], and the main benefit of frame-based approaches is eliminating the necessity to involve the tensor product, thus greatly improving the computational efficiency.
@@
Herein, we propose a local-frame-based equivariant atomistic model, named AlphaNet, for accurate yet highly efficient atomistic simulations. Building on the success of frame-based NNIPs in small atomistic systems, we introduce an additional rotary position embedding to enhance the frame transition and temporal connection for multi-scale modeling. Extensive quantitative benchmarks demonstrate that AlphaNet excels at accuracy, efficiency and scalability compared with the state-of-the-art (SOTA) NNIP models on a variety of datasets, including formate decomposition, defected graphenes, zeolites, Open Catalyst 2020 (OC20), and Matbench discovery WBM test set, which cover various types of interatomic interactions across a rich set of systems and phenomena, ranging from molecules and bulks to surfaces, from gas-phase reactions and crystal stability to surface catalysis.
@@ 2 Results
We systematically assess the capabilities of AlphaNet by examining its prediction accuracy, scalability with respect to model and dataset sizes, as well as computational efficiency in terms of inference speed and memory usage. Initially, we validate atomic-level prediction accuracy across five diverse chemical systems, simultaneously demonstrating AlphaNet’s scalability concerning dataset and system sizes. Subsequently, we quantify computational efficiency by benchmarking both the memory footprint and inference costs of the model. Comprehensive details regarding datasets, comparative methodologies, and hyperparameter configurations are provided in Appendix B.
2.1 AlphaNet achieving high accuracy across all datasets
We first consider the formate decomposition dataset, which serves as a representative example of catalytic surface reactions, specifically focuses on the dehydrogenation of formate (HCOO* → H* + CO2) on a Cu $\langle 110\rangle$ surface. For this dataset, AlphaNet achieves a mean absolute error (MAE) of 45.5 meV/Åfor force and 0.23 meV/atom for energy, compared to NequIP [5] 47.3 meV/Åand 0.50 meV/atom, respectively, as shown in Table 1. These results underscore AlphaNet’s ability to effectively capture the intricate nature of heterogeneous systems, which often involve both metallic and covalent bonding, along with complex charge transfer processes between the metal and adsorbed molecules. This high degree of accuracy highlights the model’s suitability for simulating catalytic reactions with multiple interaction types.
@@ 2.1 AlphaNet achieving high accuracy across all datasets
Defected Graphene Force MAE (meV/Å) $\downarrow$ 60.2 19.4
Energy MAE (meV/atom) $\downarrow$ 1.9 1.2
A further challenge for neural network interatomic potentials (NNIPs) is the accurate modeling of inter-layer sliding effects in layered materials such as defected graphene, where interlayer interactions are essential. On the defected graphene dataset, AlphaNet attains a force MAE of 32.0 meV/Å and an energy MAE of 1.7 meV/atom, significantly outperforming NequIP’s 60.2 meV/Å and 1.9 meV/atom, respectively. Additionally, AlphaNet successfully reproduces the binding energy profile for AB-stacked bilayer graphene, benchmarked against PBE+MBD calculations, without explicitly incorporating long-range dispersion corrections. Deviations within the shallow sliding potential energy landscape are less than 0.4 meV/atom, highlighting AlphaNet’s exceptional performance in modeling subtle interlayer forces and complex structural dynamics (see Figure 1).
@@
The third dataset evaluated is the zeolite dataset, consisting of 16 distinct zeolite types and comprising a total of 800,000 configurations. As summarized in Table 2, AlphaNet demonstrates significantly improved performance compared to Deep Pot [56] across nearly all zeolite configurations evaluated, exhibiting superior accuracy in both energy and force predictions. For example, in the ABWopt configuration, AlphaNet achieves an energy mean absolute error (MAE) of 12 meV, markedly lower than Deep Pot’s 90 meV. Similarly, AlphaNet attains a force MAE of 19 meV/Å, clearly outperforming the 90 meV/Å obtained by Deep Pot.
@@
Energy MAE $\downarrow$ 90 12 110 6.8 210 29 150 18
Force MAE $\downarrow$ 90 19 50 12 60 16 90 15
@@
Energy MAE $\downarrow$ 40 8.7 290 43 60 11 150 10
Force MAE $\downarrow$ 50 13.5 130 28 50 11 70 13
@@
Energy MAE $\downarrow$ 110 21 150 19 70 15 210 16
Force MAE $\downarrow$ 70 12 64 12 110 11 110 15
@@
Energy MAE $\downarrow$ 90 18 60 20 200 9.5 160 17
Force MAE $\downarrow$ 70 24 120 14 110 12 60 15
@@
In addressing the next challenge, we consider the Open Catalyst Project OC20 dataset [11], a comprehensive collection of crystal structures specifically curated for training neural network interatomic potential (NNIP) models in catalysis applications. Our study focuses on the OC2M subset and evaluates performance on the Structure-to-Energy-and-Force (S2EF) task. We benchmark our proposed model, AlphaNet, against established state-of-the-art methods, including EquiformerV2 [36], EScAIP [42], eSCN [41], GemNet-OC [24], SchNet [46], and DimeNet++ [22]. AlphaNet, trained on the 2M subset for 2,000,000 steps, achieves remarkable accuracy in energy prediction, yielding a mean absolute error (MAE) of 0.24 eV. This accuracy is on par with larger-scale models such as EquiformerV2 and EScAIP. In contrast, models like SchNet and DimeNet++, even when trained on the full dataset, demonstrate inferior performance, with MAE values exceeding 0.35 eV. However, AlphaNet, trained exclusively on the smaller 2M subset, does not surpass the larger-scale models (EquiformerV2 and EScAIP) in force prediction tasks. This discrepancy likely arises from the inherent conservativeness of the model architectures, which enforce consistency by computing forces strictly as negative gradients of the predicted energies ($-\nabla_{x}E(x)$).
@@
EScAIP 0.782 5.634 0.712 0.939 0.038 0.783 45M EFSD
AlphaNet-L 0.799 4.863 0.743 0.933 0.041 0.745 16.2M EFSG
AlphaNet-S*111* means not submitted yet 0.808 4.915 0.751 0.935 0.037 0.796 4.5M EFSG
Lastly, we validate the effectiveness of AlphaNet using the established Matbench discovery WBM test set [44]. AlphaNet demonstrates an excellent balance between predictive accuracy and physical consistency on the WBM benchmark (Table 4). Achieving an F1 score of 0.799 and a DAF score of 4.863, AlphaNet attains near state-of-the-art classification performance while utilizing an EFSG-type conservative force field to rigorously enforce physical constraints. Notably, despite its compact architecture (16.2M parameters), AlphaNet-L surpasses the larger eqV2 S DeNS model (31.2M parameters) in regression tasks, yielding an improved coefficient of determination ($R^{2}=0.742$) and lower prediction error (mean absolute error, $\text{MAE}=0.039$). There is also a new small size model(AlphaNet-S) with a different RBF function which performs even better. This exceptional combination of computational efficiency and physics-informed design positions AlphaNet as a powerful neural network interatomic potential (NNIP) model, ideally suited for scalable materials modeling.
2.2 AlphaNet scaling effectively with model, data and system sizes
In addition to demonstrating robust performance on large datasets and extensive systems, we systematically evaluate how model size and dataset scale influence performance using the zeolite dataset. As depicted in Figure 2, increasing the training dataset size from 50k to 800k samples notably improves prediction accuracy. Specifically, the mean absolute error (MAE) for energy predictions decreases significantly from 220 meV at 50k samples to 41 meV at 800k samples. Similarly, the force prediction MAE reduces markedly from 241 meV/Å at 5k samples to 54 meV/Å at 80k samples. Furthermore, deeper neural networks exhibit accelerated performance improvements as dataset size increases, underscoring AlphaNet’s scalability with respect to both dataset and model size.
@@ 2.2 AlphaNet scaling effectively with model, data and system sizes
2.3 AlphaNet demonstrating enhanced speed-accuracy trade-off
The primary advantage of NNIP models over density functional theory (DFT)-calculated energies and forces is their substantially faster inference speed. However, increasing the number of parameters in an NNIP model to enhance its predictive accuracy typically results in slower inference speeds. To systematically assess inference speed across various NNIPs, we measure the average forward-pass computation time using 2,000 zeolite structures, processed with a batch size of 10. Additionally, we evaluate inference performance across systems of varying sizes by systematically expanding their unit cells. While we strive to maintain a consistent parameter count across models, minor variations persist due to differing internal architectures and components, as detailed in Section B.4.
(a) Inference speed as a function of system size on the zeolite dataset.
(b) Inference speed vs performance on Matbench Discovery.
Figure 3: Comparisons of inference speed among various NNIP models. (a). The x-axis represents the number of atoms in the system, and the y-axis shows the average time required to predict energy and forces for systems in batches of 10, across a total of 200 batches on the Zeolite dataset. (b). All models are accessible and compliant with the Matbench Discovery leaderboard. The speed is calculated as the number of energy evaluation per second.
As shown in Figure 3(a), AlphaNet exhibits a clear computational advantage over other NNIP models, characterized by a notably slower increase in computational time with growing system size. This efficient scaling behavior makes AlphaNet particularly well-suited for simulations involving larger systems. Furthermore, we assess the inference speed of pre-trained models featured on the Matbench Discovery leaderboard [44]. To ensure a fair comparison, we select compliant, open-source models that are freely accessible without additional authorization. The average energy prediction speeds of these models are presented in Figure 3(b). The evaluated models include AlphaNet, ORB-v2 [39], SevenNet [40], GRACE [8], MACE [4], and CHGNet [15]. Among these, GRACE implemented using TensorFlow demonstrates significantly faster inference compared to the other models, which are built on PyTorch, and our model, AlphaNet-S, comes in the second place. In particular, AlphaNet-L exhibits an inference speed comparable to models possessing approximately ten times fewer parameters, such as SevenNet-l3i5. This finding indicates that AlphaNet maintains competitive computational efficiency despite its relatively larger size, raising an interesting prospect for investigating its potential performance with reduced parameters in future studies.